Can you help me homework ?
3. (8,1), (1,y); m = -1
I don't understand how to do this. Can you explain it to me ?

Answer by Ed Parker(21)

3. (8,1), (1,y); m = -1
I don't understand how to do this. Can you explain it to me ?
You are given a point (ordered pair) (8,1) and a slope m = -1.

So you begin by using the point-slope formula:

y - y1 = m(x - x1)  where (x1, y1) = (8,1)

Plugging in and solving for y

y - 1 = -1(x - 8)
y - 1 = -x + 8
    y = -x + 9

Now you are asked to find  y in (1,y)

That just asks, "What is y when x is 1?"

So we just plug in 1 for x and solve for y:

    y = -x + 9
    y = -1 + 9
    y = 8

So (1,y) is (1,8)

Edwin